OpenStudy (anonymous):

q varies jointly with v and s. When v=2 and s=3, q=24. Find the constant of variation, then write an equation to describe the relationship then use your equation to find q when v=4 and s=2.

OpenStudy (whpalmer4):

Do you know what joint variation means?

OpenStudy (anonymous):

yes

OpenStudy (unklerhaukus):

q varies jointly with v and s \[q\propto vs\\ q=kvs\]

OpenStudy (anonymous):

every z is the same # times x & y. (z varies jointyly with x & y. Model = kxy

OpenStudy (whpalmer4):

Right. So, what is \(k\) in this case? \[q = k * v *s\]\[24 = k*2*3\]\[k =\] Then use that value of \(k\) to evaluate \[q = k * v * s\] with \(v=4,\, s=2\)

OpenStudy (whpalmer4):

You can also shortcut the latter computation by noticing that the first time around you have k*2*3 and the second time you have k*4*2, so the new value of q will simply be the old value of q*4/3

OpenStudy (anonymous):

okay, thank you :-)